Problem 20-29 (LO. 1, 4)

Pink Corporation acquired land and securities in a § 351 tax-free exchange. On the date of the transfer, the land had a basis of $720,000 and a fair market value of $630,000, and the securities had a basis of $110,000 and a fair market value of $250,000. Pink Corporation has two shareholders, Maria and Paul, who are unrelated. Maria owns 85% of the stock in The Corporation, and Paul owns 15%. Pink adopts a plan of liquidation in the current year. On the date of the liquidation, the land's fair market value has decreased to $500,000. What is the effect of each of the following on Pink Corporation?

If an amount is zero, enter "0".

a. If Pink Corporation distributes all the land to Maria, the corporation has a realized loss  $ of which
$ will be recognized.

b. Assume there is no business reason for the transfer of the property.

If all the land is distributed to Paul, then the property had a built-in loss  of $ which is disallowed . Therefore, Pink Corporation will recognize a loss  of $.

c. Assume there is no business reason for the transfer of the property. If the land is distributed 85% to Maria and 15% to Paul, Pink Corporation has a disallowed loss of $ due to the distribution of disqualified property. In addition, $ is disallowed due to the built-in loss rules. As a result, only $ of the loss will be recognized

d. Assume there is no business reason for the transfer of the property. If the land is distributed 50% to Maria and 50% to Paul, the
$ loss on the distribution to Maria, is disallowed.  Of the $ loss on the distribution to Paul,
$ will be disallowed.

e. If the land is sold and the proceeds of $500,000 are distributed proportionately to Maria and to Paul but Pink Corporation cannot show a business purpose for the transfer, the built-in loss  of $ will be disallowed.
Answer #1

ANSWER

a).

realized loss = 500,000 - 720,000 = 220,000

b).

Built in loss = 90,000 [ 630,000 - 720,000]

c).

Loss on distribution = 220,000 *85% = 187,000

Built in loss = 90,000 *15% = 13500

Loss recognized = 220,000 - 187,000 - 13500 = 19500

d).

Loss on distribution to Maria = 110,000 [ 220,000*50%]

Loss on distribution to Paul = 110,000

Loss disallowed = 110,000

e).

Built in loss =$90,000

__________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU*************